THE PERMANENT ELECTRIC DIPOLE MOMENTS OF THE X^{3}Δ, E^{3}Π, A^{3}Φ AND B^{3}Π STATES OF TITANIUM MONOXIDE, TiO

\underline{TIMOTHY C. STEIMLE} and WILTON L. VIRGO, Department of Chemistry and Biochemistry, Arizona State University, Tempe, Arizona 85287-1604; JAMES T. MUCKERMAN, Chemistry Department, Brookhaven National Laboratory, Upton, NY 11973-5000.

The optical Stark spectrum of the origin bands of the E^{3}Π_{0} − X^{3}Δ_{1}, A^{3}Φ_{2} − X^{3}Δ_{1}, and B^{3}Π_{0} − X^{3}Δ_{1} band systems of titanium monoxide, TiO, were analyzed to produce permanent electric dipole moments of 3.34(1)D, 3.2(4)D, 4.89(5)D and 4.9(2)D for the X^{3}Δ_{1}, E^{3}Π_{0}, A^{3}Φ_{2} and B^{3}Π_{0} states, respectively. The observations are compared with a simple molecular orbital description for the low-lying states and electronic structure calculations.